Find all points on the y-axis that are 6 units from P(4,1)
Enter answers as two ordered pairs.
Round answers to nearest hundreth.

Answer by Alan3354(69443)

Find all points on the y-axis that are 6 units from P(4,1)
Enter answers as two ordered pairs.
Round answers to nearest hundreth.
----------------
The 2 points will be the intersection of a circle of radius 6 centered at (4,1) and the y-axis
The circle is:
(x-4)^2 + (y-1)^2 = 36
The y-axis is x = 0
(-4)^2 + (y-1)^2 = 36
y^2 - 2y + 1 = 20
y^2 - 2y - 19 = 0
----------------
y = 1 + 2sqrt(5) =~ 5.47 b--> (0,5.47)
y = 1 - 2sqrt(5) =~ -3.47--> (0,-3.47)